OpenStudy (anonymous):

can some one show me this. I just don't want the answer. m=5/6, (2,-4). The equation of the line is y=

OpenStudy (anonymous):

Do you all need more info on this? I would like someone to show me how to do this please.

OpenStudy (anonymous):

You know the equation of the line is y=mx+b, correct?

OpenStudy (anonymous):

If so your problem gives you your slope (m), and a point it will pass through (2, -4). So plug those numbers in, and solve for b, because that is the number we need to make the equation of a line. So -4=5/6(2)+b. I'll leave it to you to solve for b. Once you find b, then you can write y=(5/6)x+(whatever the number you found b to be)